Part A How many mL of H2O needed in these dilutions? Vol of 1mM HCI (ml) Vol H20 needed (ml) Dilution 1 to 2 1 to 10 1 to 5 1 to 4 1 to 12 0.5 0.3 0.8 Part B You are given 50 mL solution of NaCl with a concentration of 2M. Find the dilutions needed and the volumes needed of NaCI with the info given NaCl concentration Final Required concentration Dilution 2M 2M 2M 2M 2M Final Volume (ml) Vol of NaCl (ml) 100 40 10 1 to 1 to 100 mM 200mM 500 uM 1000mM 500mM 1 to 30

Show calculations for finding these ratios to create dilutions given the concentrations.
Answer #1

Ans. # Part A: Dilution factor = Vol. of stock soln. / Final vol. of soln. made upto

#1. Final vol. of soln. made up to = Vol. of stock soln. / Dilution factor

                                                            = 1 mL / (1/2)

                                                            = 2 mL

Now,

            Required Vol. of H2O = Final vol. of soln. made upto – Vol. of stock soln.

                                                = 2.0 mL – 1.0 L = 1.0 mL

# Rest of the calculations are summarized in excels sheet.

Vol. of Stock HCl Final vol., mL Req. vol. Of H20 mL, V1 V2 (V/DF) mL V2- V1) Dilution, DF 1 to2 1 to 10 1.00 1.00 0.50 0.30 0.80 2.00 10.00 2.50 1.20 9.60 1.00 9.00 2.00 0.90 8.80 1 to <4 1 to 12

# Part B: #1. Using C1V1 (stock soln.) = C2V2 (final soln)

            V1 = (100 mM x 100 mL) / 2 M = (0.10 M x 100 mL) / 2 M = 5.0 mL

Now,

            Dilution factor = 5.0 mL / 100 mL = 1 / 20

# Rest of the calculations are summarized in excel sheet.

Vol. of Stock HCl Final vol., mL Req. vol. Of H2O V2 (= V/ DF) mL (= V2-V1) Dilution, DF 1 to 2 1 to 10 mL, V1 1.00 2.00 0.50 0.30 0.80 2.00 20.00 2.50 1.20 9.60 1.00 18.00 2.00 0.90 8.80 1 to 4 1 to 12